The radius of a circular flower garden is 4 meters. How many meters of edging will be needed to surround the garden?
A) 7.14 meters B) 12.56 meters C)25.12 meters D)50.27 meters E)3.14 meters

Answer by richwmiller(17219)

c=2 pi r
c=2*3.14*4
c=8*3.14
We can approximate 8*3=24 the only answer even close is 25.12
since "none of the above' is not a choice it must be 25.12.
and yes 3.14*8=25.12
if you didn't know that 8*14 is 112 you should know that 8*7=56 and 2*56=112